OpenStudy (anonymous):

What is the standard form of the equation of a circle with its center at (2, -3) and passing through the point (-2, 0)?

OpenStudy (anonymous):



OpenStudy (anonymous):

the distance between (-2,0) and (2,-3) will give us the radius

OpenStudy (anonymous):

\[\sqrt{(-3-0)^{2}+(2-(-2)^{2}}\]

OpenStudy (anonymous):

radius is 5

OpenStudy (anonymous):

i tried doing the problem but i got it wrong . i put the radies as 25 @strenesmee

OpenStudy (anonymous):

(x-2)^2+(y-(-3))^2=25 x^2-4x+4+y^2+6y+9=25 x^2-4x+y^2+6y-12=0

OpenStudy (anonymous):

what about this ? im not very good at geomtry & im trying but its kind of confusing , what about this A circle is described by the equation (x−12)2+(y−32)2=49. What are the cordinates for the center of the circle and the length of the radius?

OpenStudy (anonymous):

@strenesmee

OpenStudy (anonymous):

the equation for a circle is like this (x-a)^2+(y-b)^2=r^2 here a stands for x axis of the center and b stands for y axis of the center and r is the radius of the circle can you guess the answer??

OpenStudy (anonymous):

(-1/2,-3/2), 2/3units (-1/2,-3/2), 4/9units (1/2,3/2), 4/9units (1/2,3/2), 2/3units these are the answer choices . i cant really take a wild guess

OpenStudy (anonymous):

first try to get an equation like that (x-a)^2+(y-b)^2=r^2